# AnchorCrafter: Animate CyberAnchors Saling Your Products via Human-Object Interacting Video Generation

### v1

@@ 5.4 User Study
We conducted a user preference evaluation with 50 participants, who rated 30 videos from each method based on five criteria: appearance preservation for humans and objects, motion accuracy for humans and objects, and overall quality. Appearance preservation assessed alignment with the reference image, while motion accuracy focused on adherence to pose conditions, particularly for hands and objects. Each criterion was rated on a five-point scale, with five being the highest score. As shown in Table 2, our model consistently outperformed others across all criteria.
@@
AnyV2V [19] 1.60 1.86 1.06 1.40 1.06
MimicMotion+AnyDoor [6] 3.13 2.33 2.53 2.06 2.13
AnimateAnyone [14] 1.81 2.13 1.61 1.72 1.95
MimicMotion [41] 3.93 3.20 4.06 2.80 2.80
Ours 4.00 4.26 4.53 4.60 4.47
Table 2: User study scores. The rating score is on a scale from one to five, where five is the highest score, and one is the lowest.
5.5 Limitation
Figure 7: Limitions.
AnchorCrafter is primarily evaluated on rigid objects and doesn’t perform well on transparent or non-rigid bodies. In Fig. 7, it fails with mirror penetration. In the future, we will investigate the challenges associated with transparent and non-rigid objects.
@@ 6 Conclusion
We present AnchorCrafter, a novel diffusion-based system for anchor-style product promotion video generation, incorporating human-object interaction into pose-guided human video generation. By introducing HOI-appearance perception and HOI-motion injection, our system addresses key challenges in object motion guidance, appearance preservation, and complex human-object interactions. We also propose an HOI-region reweighting loss to improve object details during training. Extensive experiments show that AnchorCrafter outperforms existing methods, achieving superior object appearance preservation and shape awareness while ensuring high-quality video generation with consistent human appearance and motion.

### v2

@@ 5.4 User Study
TABLE II: User study scores. The rating score is on a scale from one to five, where five is the highest score and one is the lowest.
@@
MimicMotion [7] 3.52 1.84 3.68 2.08 3.08
StableAnimator  [9] 3.02 2.18 3.64 1.96 2.96
UniAnimate-DiT  [23] 4.02 1.28 3.34 1.56 3.14
VACE  [24] 2.28 1.92 2.42 2.16 2.62
Ours 4.12 4.68 4.58 4.82 4.64
We conducted a user preference evaluation comparing AnchorCrafter with four state-of-the-art methods. Each participant reviewed 30 randomly selected videos for each method, scoring them based on five criteria:
  • Appearance Preservation (Human): Evaluates the consistency between humans and the reference images.
  • Appearance Preservation (Object): Focuses on maintaining object details such as texture and shape.
  • Motion Accuracy (Human): Assesses the precision of human movements, particularly poses and gestures.
  • Motion Accuracy (Object): Ensures that objects’ motions align with intended physics and interactions.
  • Overall Quality: Provides a comprehensive evaluation of visual appeal, coherence, and stability.
Each criterion was rated on a five-point scale, with five being the highest score. Before the evaluation, participants were given detailed instructions, reference videos, and corresponding conditions to ensure fairness and clarity throughout the process. A total of 50 participants were invited, including 25 males and 25 females, achieving balanced gender representation. The participants consisted of 30 computer vision researchers, 10 artists, and 10 individuals from diverse backgrounds. This diverse group provided comprehensive feedback from various perspectives. As shown in Table II, our approach consistently ranked the highest across all five criteria, excelling particularly in motion accuracy and appearance diversity. This demonstrates its strength in generating realistic and stable human-object interaction videos.
Figure 9: Anchorcrafter is able to perform the action of ”opening the headphone case”, demonstrating the adaptability and interactive ability of our model in complex scenarios.
5.5 Discussions
5.5.1 Generalization in Multi-View Tasks
Human-object interaction tasks frequently require displaying objects from diverse perspectives. To this end, the proposed method utilizes multiple images to capture intricate details from various viewpoints. For simplicity, the experiment emphasizes the front view and two side views of the object. In practice, these views can be substituted with other perspectives during fine-tuning, such as the object’s rear or even its interior. As shown in Fig. 9, leveraging depth perception and multi-view images, we successfully achieve the complex task of ”opening the headphone case”. This approach effectively captures detailed features from various angles of the headphones, showcasing our model’s adaptability and interaction capabilities in intricate scenarios.
5.5.2 Generalization Across Shape-Similar Objects
Figure 10: Generalization over similar object shapes. Driving iPhone interaction videos with another mobile phone.
During inference, generate interaction videos with a specific object but don’t have access to it, can use a similarly shaped object instead. By capturing footage with a similar object, we can extract human poses, 3D hand mesh, and object depth as conditioning inputs. Our model exhibits a certain degree of generalization capability, enabling accurate inference even with similar-shaped objects. As shown in Fig. 10, we utilized videos recorded with different phone models to drive video generation for an iPhone. For objects with similar shapes, template videos can be captured to guide multiple objects.
Figure 11: Limitations. Handling transparent objects remains challenging, and erroneous conditions introduce interference.
5.6 Limitations
AnchorCrafter is primarily evaluated on rigid objects and exhibits suboptimal performance when handling transparent objects. Additionally, inaccuracies in preprocessing models such as DWPose introduce errors into the generated results. As shown in Fig. 11 (left), the model fails in cases of mirror penetration. In Fig. 11 (right), the 3D hand mesh provides accurate hand guidance, whereas DWPose incorrectly assigns the left hand, leading to erroneous generation results. In future work, we will conduct a more in-depth exploration of the challenges associated with transparent and non-rigid objects.
@@ 6 Conclusion
We present AnchorCrafter, a novel diffusion-based system for anchor-style product promotion video generation that incorporates human-object interaction into pose-guided human video generation. Our system addresses key challenges in object motion guidance, appearance preservation, and complex human-object interactions by introducing HOI-appearance perception and HOI-motion injection. We also propose an HOI-region reweighting loss to improve object details during training. Extensive experiments show that AnchorCrafter outperforms existing methods, achieving superior object appearance preservation and shape awareness while ensuring high-quality video generation with consistent human appearance and motion.